OpenStudy (anonymous):

Solve the equation n/4 = 8

OpenStudy (saifoo.khan):

multiply 4 on both sides, \[n = 4 * 8\]\[n = 32\]

OpenStudy (anonymous):

32

OpenStudy (karatechopper):

n/4=8 n/4 x 4=8 x 4 n=32